OpenStudy (anonymous):

Solve for x 3x^2-6x=-3

OpenStudy (anonymous):

quadratic equation here 3x^2-6x+3=0 factorise common factors 3(x^2-2x+1) = 0 x^2-2x+1=0 factorise this perfect square (x-1)^2 = 0 therefore, x - 1 = 0 x = 1

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

anytime ;)